OpenStudy (anonymous):

If the quiz consists of 3 questions, question 1 has 3 possible answers, question 2 has 4 possible answers, and question 3 has 5 possible answers, find the probability that Erin gets one or more correct answers.

OpenStudy (anonymous):

easy method is to compute the probability that "erin" gets none then subtract that number from one

OpenStudy (anonymous):

Let X = the number of correct answers. We can write X = X1 + X2 + X3, where Xi = the number of correct answers on question i. (Note that the only possible values of Xi are 0 and 1, with 0 representing an incorrect answer and 1 a correct answer.) The probability of at least one correct answer is P(X ≥ 1) = 1 − P(X = 0) = 1 − [P(X1 = 0) (X2 = 0) P(X3 = 0)] (since the Xi are independent) =

OpenStudy (anonymous):

\[1 - (2/3) (3/4) (4/5) = 1- 24/60 = 0.6\]

OpenStudy (anonymous):

And if the question has n possible answers, the probability of guessing it right is 1/n and of guessing it wrong is (n-1)/n

OpenStudy (anonymous):

1 - 0.4 = .60 ^^ theres your answer